# SReferences

### v1

@@ II RUS fusion with re-initialization of emitters
For RUS logical fusions, when a photon is lost during a physical fusion, the $XX$ outcome is irreversibly lost thus $\overline{XX}$ cannot be recovered. On the other hand, the $\overline{ZZ}$ outcome can still be obtained by retrieving $ZZ$ in the next fusion attempt through biasing [10], due to re-emission of the photons followed by single-qubit measurement in the Pauli-$Z$ basis. Crucially, when the $ZZ$ outcome is retrieved, the two emitter spins are indirectly measured in the Pauli-$Z$ basis, since the measured photon and emitter share the stabilizer $ZZ$. This implies that once the logical fusion results in $\overline{ZZ}$ outcome only, the emitter spins become disconnected from the original graph with their states effectively re-initialized.

### v2

@@ III RUS fusion with re-initialization of emitters
For RUS logical fusions, when a photon is lost during a physical fusion, the $XX$ outcome is irreversibly lost thus $\overline{XX}$ cannot be recovered. On the other hand, the $\overline{ZZ}$ outcome can still be obtained by retrieving $ZZ$ in the next fusion attempt (Sec. VI.2), due to re-emission of the photons followed by single-qubit measurement in the Pauli-$Z$ basis. Crucially, when the $ZZ$ outcome is retrieved, the two emitter spins are indirectly measured in the Pauli-$Z$ basis, since the measured photon and emitter share the stabilizer $ZZ$. This implies that once the logical fusion results in $\overline{ZZ}$ outcome only, the emitter spins become disconnected from the original graph with their states effectively re-initialized.